7.) A screw was detached from a hot air balloon that is rising at 3.5 m/s at an altitude of 8.3 m. (a) What is the initial velocity of the screw? (b) How long will it take for the screw to reach the ground?
1
Expert's answer
2019-08-22T09:27:54-0400

First a reference system is established.


Where semi positive axis points up + and


Question (a)


The initial velocity of the screw is equal to the initial velocity of the balloon at the time the screw separates, therefore


"\boxed{V_{oy}=3.5\frac{m}{s}}" .


Question (b)


The screw position is given by:


"y_{f}=y_{o}+V_{oy}*t+\frac{1}{2}a_{y}t^{2}"


Where.

  • Final position (when it reaches the ground)

"y_{o}=8.5m"

  • Initial position

"y_{f}=0m"

  • Initial velocity

"V_{0y}=3.5\frac{m}{s}"

  • Gravity acceleration

"a=-9.8\frac{m}{s^{2}}"


Numerically evaluating and organizing

 

A second grade equation.


"0=8.5+3.5t-\frac{9.8}{2}t^{2}"


"-4.9t^{2}+3.5t+8.5=0-->ax^{2}+bx+c=0"


Coefficients "a=-4.9;b=3.5;c=8..5"


solving numerically


"t=\frac{-b\pm \sqrt{b^{2}-4*a*c}}{2*a}"


"t=\frac{-(3.5)\pm \sqrt{(3.5)^{2}-4*-4.9*8.5}}{2*-4.9}"


  • Solution 1

"t=\frac{-(3.5)+ \sqrt{(3.5)^{2}-4*-4.9*8.5}}{2*-4.9}=-1.00s"


  • Solution 2

"t=\frac{-(3.5)- \sqrt{(3.5)^{2}-4*-4.9*8.5}}{2*-4.9}=1.72s"


The time considered is positive.


The time it takes to get to the ground is:


"\boxed{t=1.72s}"